# Bayesian Multifractal Image Segmentation

### v1

@@ Abstract
Multifractal analysis (MFA) provides a framework for the global characterization of image textures by describing the spatial fluctuations of their local regularity based on the multifractal spectrum. Several works have shown the interest of using MFA for the description of homogeneous textures in images. Nevertheless, natural images can be composed of several textures and, in turn, multifractal properties associated with those textures. This paper introduces a Bayesian multifractal segmentation method to model and segment multifractal textures by jointly estimating the multifractal parameters and labels on images. For this, a computationally and statistically efficient multifractal parameter estimation model for wavelet leaders is firstly developed, defining different multifractality parameters to different regions of an image. Then, a multiscale Potts Markov random field is introduced as a prior to model the inherent spatial and scale correlations between the labels of the wavelet leaders. A Gibbs sampling methodology is employed to draw samples from the posterior distribution of the parameters. Numerical experiments are conducted on synthetic multifractal images to evaluate the performance of the proposed segmentation approach. The proposed method achieves superior performance compared to traditional unsupervised segmentation techniques as well as modern deep learning-based approaches, showing its effectiveness for multifractal image segmentation.
@@ 1 Introduction
Scaling, or scale invariance, is a modeling paradigm in which a given object structure or statistical property remains unchanged regardless of the scale at which it is analyzed [1], and can be exploited using self-similar or multifractal models [2]. The scaling properties of texture images have been studied via the multifractal analysis (MFA) tool that allows complex patterns (local transient regularity fluctuations) to be described and modeled across different scales [3, 4]. MFA has been exploited in several applications of different natures, including network traffic analysis [5], texture classification [6, 7, 8], texture characterization in general [9] and art analysis [1], among others. Current state-of-the-art MFA tools rely on the use of dedicated multiresolution coefficients, the wavelet leaders [2, 7, 10].
@@
On the other hand, texture segmentation is a long-standing problem in image processing that has received significant research efforts in the past and continues to attract considerable attention in the scientific community [11, 12, 13]. Several methods have relied on geometrical or statistical features using Markov random fields [14, 15], while others rely on pixel intensities to group different regions [16, 17]. Other works have exploited the scale invariance property for image segmentation [18, 19]. As an example, self-similarity exponents and local regularity have been used to segment textures on images [18]. Furthermore, the multifractal spectrum can also be used to determine features that can be input to the $k$-means clustering algorithm [19]. Works such as [20] have also used fractal information to extract features. In this line, authors in [4] provide a methodology for modeling the multifractality of multivariate images, yielding an operational tool for the multifractal analysis of multidimensional images as well as images with several multifractal textures. However, the method is not intended to assess multifractality at the pixel level or for segmentation purposes, which is a key limitation of current multifractal-based approaches. Indeed, characterizing point-wise regularity fluctuations requires to analyze several scales, tied to space averages [21], thus yielding a paradox and an important difficulty for pixel-level analysis and segmentation.
@@
Multi-resolution coefficients associated with an image have been used in several works to perform image segmentation. Works such as [22] and [23] have exploited, for instance, the wavelet transform to compute multi-resolution coefficients of an image that are combined from coarse to fine scales for image segmentation. Further works such as [24, 25] have proposed multiscale Bayesian techniques in which a contextual behavior from a coarse scale (a low-resolution scale) is employed up to the finest scale (highest resolution) to provide an image segmentation. Multiscale segmentation has also been performed using hidden Markov trees [22, 26] or neural networks [27]. The main limitation of these approaches is that they rely on training data to describe the texture parameters of the images to be segmented. Additionally, even though there exist works in the literature that attempt to consider coefficients at multiple scales for image segmentation, up to our knowledge, there is no method in the state of the art that exploits these scales to segment images at the pixel level based on a multifractal model.
@@
This work introduces a multifractal image segmentation approach for the joint parameter and label estimation of heterogeneous images via a multiscale decomposition and a Bayesian framework. The proposed segmentation approach is fully unsupervised and is conducted in the wavelet leader domain, in which the scale-free textures are parameterized by the multifractality parameter. The contributions of this paper are the following:
@@
    First, a computationally efficient Fourier-based Whittle approximation for the multifractality parameter estimation on heterogeneous multifractal images is introduced. This contribution relies on the efficient model developed in [4] that has been extended and adapted, here, to irregular lattices for modeling the different piecewise textures of an image.
@@
    Third, a multiscale Potts Markov random field that considers the intra-scale and inter-scale dependencies in the wavelet leader domain is proposed. This random field uses a multiscale graph structure that exploits spatial and scale image correlations. Moreover, the granularity coefficients of the proposed multiscale Potts-Markov random field are efficiently and jointly sampled with the labels.
@@
    Fourth, we conduct extensive numerical experiments and comparisons with state-of-the-art image segmentation methods for simulated and real-world images, showing the interest of the proposed framework and its superior performance by a margin.
@@ 2.1 Local Regularity and Multifractal Spectrum
Multifractal analysis (MFA) enables the study of local regularity of functions. To that end, it makes use of a quantity that measures the so-called Hölder exponent $h(\textbf{t})$ at a location t, and a global description of its spatial fluctuations, the multifractal spectrum $\mathcal{D}(h)$ [28].
@@
Here, large (small) values of $h(\textbf{t}_{0})$ correspond to locations where $\mathbf{F}$ is locally smooth (irregular), respectively [18].
@@
The estimation of $\mathcal{D}(h)$ is the main target of MFA. However, it is unfeasible using its formal definition (1). Indeed, the estimation relies on a multiscale decomposition and the so-called multifractal formalism (for more details, see [2, 4, 10]). The current state-of-the-art multifractal formalism is based on the wavelet leaders of the two-dimensional (2D) discrete wavelet transform (DWT) coefficients of the given image and is recalled next.
@@ 2.2 Multifractal Formalism based on Wavelet Leaders
Wavelet Leaders.  The 2D orthonormal DWT of $\textbf{F}\in\mathbb{R}^{N\times N}$ is obtained by considering four 2D filters $\phi^{(m)}(\textbf{n})$, $m=0,...,3$, defined as the tensor products of a low-pass filter $\phi(n)$ (scaling function) and a high-pass filter $\psi(n)$ (wavelet function characterized by $N_{\psi}>0$ vanishing moments). Precisely, the four 2D filters are written as [29, 4]
@@
where $\psi^{(0)}(\textbf{n})$ is the 2D low-pass filter and $\{\psi^{(m)}(\textbf{n})\}_{m=1}^{3}$ are the 2D high-pass filters. Let $D^{(m)}_{\textbf{F}}(j,\mathbf{n})$ denote the 2D DWT of F, with $D^{(0)}_{\textbf{F}}(0,\mathbf{n})\triangleq\textbf{F}$. The DWT coefficients $\big{\{}D_{\textbf{F}}^{(m)}(j,\mathbf{n})\big{\}}_{m=1}^{3}$, and the approximation coefficients $D_{\textbf{F}}^{(0)}(j,\mathbf{n})$, for $j\geq 1$, are obtained by iteratively convolving $\psi^{(m)}$ with the coefficients $D_{\textbf{F}}^{(0)}(j-1,\cdot)$, followed by decimation [4]. For MFA purposes, the 2D DWT coefficients are normalized as $d_{\textbf{F}}^{(m)}(j,\mathbf{n})\triangleq 2^{-j}D_{\textbf{F}}^{(m)}(j,\mathbf{n})$, for $m=1,2,3$.
The wavelet leaders are defined as the largest DWT coefficient magnitudes within the spatial neighborhood $3\lambda_{j,\textbf{n}}$ over all finer scales as
$L_{{j,\textbf{n}}}=\underset{{\scriptscriptstyle\begin{matrix}m\in\{1,2,3\}\\ \lambda^{\prime}\subset 3\lambda_{j,\textbf{n}}\end{matrix}}}{\textrm{sup}}\big{|}{d}^{(m)}_{\lambda^{\prime}}\big{|},$ (2)
where $\lambda_{j,\textbf{n}}$ is the dyadic cube of side length $2^{j}$ at position $2^{j}\textbf{n}$, i.e., $\lambda_{j,\textbf{n}}=\{[2^{j}n_{1},2^{j}(n_{1}+1)),[2^{j}n_{2},2^{j}(n_{2}+1))\}$ [30, 31], and $3\lambda_{j,\textbf{n}}=\bigcup_{i_{1},i_{2}\in\{-1,0,1\}}\lambda_{j,(n_{1}+i_{1},n_{2}+i_{2})}$ [4, 29].
It has been shown that the coefficients $c_{p}$ are tied to the wavelet leaders $L_{j,\mathbf{n}}$ via the key relation [29]
@@
for $c_{2}<0$, see [4, 29] for further details. These so-called log-cumulants provide a relevant summary of $\mathcal{D}(h)$, e.g., $c_{1}$ is the mode of $\mathcal{D}(h)$ that can be interpreted as the average smoothness of the image $\mathbf{F}$ and is related to the Hurst parameter [32], $c_{2}$ is related to the width of $\mathcal{D}(h)$ and captures the degree of fluctuations of the local regularity of $\mathbf{F}$ and is the most important feature of multifractal functions, termed multifractality parameter .
@@
where $\widehat{\textrm{Var}}(\cdot)$ denotes the sample variance and $w_{j}$ are suitable regression weights [2, 3, 33]. The estimator (5) has low complexity but poor performance (large variance). To overcome this limitation, a Fourier-based Bayesian model was introduced for homogeneous images [4]. It is well suited for the design of joint priors and briefly recalled next.
@@ 2.3 Estimation of $c_{2}$ using a Fourier-based Bayesian model
Let ${\ell}_{j,\textbf{n}}=\ln(L_{{j,\textbf{n}}})$ denote the centered logarithms of the wavelet leader coefficients, referred to as log-leaders, at scale $j$ and at position n. The vector of log-leaders, $\boldsymbol{\ell}_{{j}}\in\mathbb{R}^{N_{j}^{2}}$, is obtained by vectorization using the lexicographic order, where $N_{j}=N/2^{j}$. Note that $j\in\llbracket j_{1},j_{2}\rrbracket$, $1\leq j_{1}\leq...\leq j_{2}\leq J$, in which $j=1$ and $j=J$ index the finest and the coarsest scales, respectively, where $\llbracket a,b\rrbracket$ denotes the set of integers ranging from $a$ to $b$. Further, let $\boldsymbol{\ell}\triangleq[\boldsymbol{\ell}_{j_{1}}^{T},...,\boldsymbol{\ell}_{j_{2}}^{T}]^{T}$ be the vector obtained by concatenation of the log-leaders at scales $j_{1},...,j_{2}$. According to [4, 29], the distribution of the log-leaders $\boldsymbol{\ell}_{j}$ can be well-approximated by a multivariate Gaussian distribution with density given by:
$p(\boldsymbol{\ell}_{j}|\theta_{1},\theta_{2})\propto{{\big{|}\,\,\mathbf{\Sigma}_{j}(\boldsymbol{\theta})\big{|}^{-1/2}}}{\exp\Big{(}-\frac{1}{2}\boldsymbol{\ell}_{j}^{T}\mathbf{\Sigma}_{j}(\boldsymbol{\theta},\left\lVert\Delta\textbf{n}\right\rVert)^{-1}\boldsymbol{\ell}_{j}\Big{)}},$ (6)
and, assuming independence between scales,
@@
[…]
@@
where $r\triangleq\left\lVert\Delta\textbf{n}\right\rVert$, $g_{1}$ and $g_{2}$ are functions defined as $g_{1}(j,r)=\max\{0,-\ln\big{(}(r+1)/(r_{j}+1)\big{)}\}$ and $g_{2}(j,r)=\max\{0,-\ln(r+1)/\ln 4\}$ with $r_{j}=\lfloor N_{j}/4\rfloor$ [34].
The likelihood (7) is problematic for direct use in a numerical algorithm but can be efficiently approximated in the spectral domain by formulating an equivalent Fourier-domain data augmented model introduced in [4]. This statistical model is based on a Whittle approximation and data augmentation allowing the suitable design of conjugate priors for the parameters ${\theta}_{1},{\theta}_{2}$ [4, 29]. More precisely, (7) is approximated as
@@
$p\big{(}\textbf{x},\boldsymbol{\mu}|$ ${\theta}_{1},{\theta}_{2}\big{)}\propto{\theta}_{1}^{-S}\exp\bigg{(}-\frac{1}{{\theta}_{1}}(\textbf{x}-\boldsymbol{\mu})^{H}G^{-1}_{1,s}(\textbf{x}-\boldsymbol{\mu})\bigg{)}$
×$\times{\theta}_{2}^{-S}\exp\bigg{(}-\frac{1}{{\theta}_{2}}\boldsymbol{\mu}^{H}G^{-1}_{2,s}\boldsymbol{\mu}\bigg{)},$ (10)
where $\boldsymbol{\mu}$ is a latent vector and $\textbf{x}=[\textbf{x}_{1}^{T},...,\textbf{x}_{S}^{T}]^{T}$ is a subset of Fourier coefficients of $\boldsymbol{\ell}$ [4, 34],
@@
where $s\in\big{\{}(j,\textbf{m}):j=\llbracket j_{1},j_{2}\rrbracket,\textbf{m}\in\Omega^{\dagger}_{j}\big{\}}$, $\Omega^{\dagger}_{j}$ indexes low frequencies (see [4, 34] for further details), and $\boldsymbol{\omega}_{s}=2\pi\mathbf{m}/\sqrt{N_{j}}$ is the given Fourier frequency, with $(j,\textbf{m})\in s$. Moreover,
${G}_{i,s}\triangleq\sum_{\textbf{n}\in\llbracket-{N_{j}},{N_{j}}\rrbracket^{2}}g_{i}(j,\left\lVert\textbf{n}\right\rVert)\exp{(-i\,\textbf{n}^{T}\boldsymbol{\omega}_{\mathbf{s}})}.$ (12)
for $i=1,2$ are real and strictly positive matrices defined on $s$. The corresponding spectral covariance of the asymptotic approximation (2.3) is expressed as
@@
which leads to standard conditional distributions and fast estimation when used in a Bayesian model [4, 34].

### v2

@@ Abstract
Multifractal analysis (MFA) provides a framework for the global characterization of image textures by describing the spatial fluctuations of their local regularity based on the multifractal spectrum. Several works have shown the interest of using MFA for the description of homogeneous textures in images. Nevertheless, natural images can be composed of several textures and, in turn, multifractal properties associated with those textures. This paper introduces an unsupervised Bayesian multifractal segmentation method to model and segment multifractal textures by jointly estimating the multifractal parameters and labels on images, at the pixel-level. For this, a computationally and statistically efficient multifractal parameter estimation model for wavelet leaders is firstly developed, defining different multifractality parameters for different regions of an image. Then, a multiscale Potts Markov random field is introduced as a prior to model the inherent spatial and scale correlations (referred to as cross-scale correlations) between the labels of the wavelet leaders. A Gibbs sampling methodology is finally used to draw samples from the posterior distribution of the unknown model parameters. Numerical experiments are conducted on synthetic multifractal images to evaluate the performance of the proposed segmentation approach. The proposed method achieves superior performance compared to traditional unsupervised segmentation techniques as well as modern deep learning-based approaches, showing its effectiveness for multifractal image segmentation.
@@ 1 Introduction
Scaling, or scale invariance, is a modeling paradigm in which a given object structure or statistical property remains unchanged regardless of the scale at which it is analyzed [abry2013van], and can be exploited using self-similar or multifractal models [wendt2007bootstrap]. The scaling properties of texture images have been studied via the multifractal analysis (MFA) that allows complex patterns (local transient regularity fluctuations) to be described and modeled across different scales [wendt2007multifractality, wendt2018multifractal]. MFA has been exploited in several applications of different natures, including network traffic analysis [abry1998wavelet], texture classification [wendt2009wavelet, Wendt20091100, Xu2010], texture characterization [schmitt2001multifractal] and art analysis [abry2013van]. Current state-of-the-art MFA tools rely on the use of dedicated multiresolution coefficients, the wavelet leaders [wendt2007bootstrap, Wendt20091100, jaffard2004wavelet].
@@
On the other hand, texture segmentation is a long-standing problem in image processing that has received significant research efforts in the past and continues to attract considerable attention in the scientific community [jain1991unsupervised, kirillov2023segment, he2017mask]. Several methods have relied on geometrical or statistical features using Markov random fields [manjunath1991unsupervised, halimi2015unsupervised], while others rely on pixel intensities to group different regions [yuan2015factorization, haralick1979statistical]. Other works have exploited the scale invariance property for image segmentation [pustelnik2016combining, abadi2006texture]. As an example, self-similarity exponents and local regularity have been used to segment image textures [pustelnik2016combining]. Works such as [chaudhuri1995texture] have also used fractal information to extract features. The literature reports some MFA-based image segmentation methods that rely on the multifractal spectrum [abadi2006texture, san2017application, zhu2019ship] (see definition in Section II-A and see a review of the methods in [lopes2009fractal]). For instance, the multifractal spectrum was used in [san2017application] to segment water bodies in synthetic aperture radar (SAR) images. However, the strategy investigated in [san2017application] is computationally demanding and relies on an empirical threshold for region delineation. Segmentation approaches based on multifractal descriptors have also been studied [xia2006morphology, vehel1994multifractal]. The state-of-the-art MFA formalism (in terms of precision and computational efficiency) is based on wavelet leaders [jaffard2006wavelet, wendt2018multifractal, combrexelle2015bayesian]. In this line, authors in [wendt2018multifractal] provide a methodology for modeling the multifractality of multivariate images, yielding an operational tool for the multifractal analysis of multidimensional images as well as images with several multifractal textures. However, the method is not intended to assess multifractality at the pixel level or for segmentation purposes, which is a key limitation of current multifractal-based approaches. Indeed, characterizing point-wise regularity fluctuations requires to analyze several scales, tied to space averages [wendt2022multifractal], thus yielding a paradox and an important difficulty for pixel-level analysis and segmentation.
@@
Multi-resolution coefficients associated with an image have been used in several works to perform image segmentation. The wavelet transform was used in [choi2001multiscale] and [brault2005unsupervised] to compute multi-resolution coefficients of an image that are combined from coarse to fine scales for image segmentation. Further works such as [bouman1994multiscale, fan2001joint] have proposed multiscale Bayesian techniques in which a contextual behavior from a coarse scale (low-resolution scale) is employed up to the finest scale (highest resolution) to provide an image segmentation. Multiscale segmentation has also been performed using hidden Markov trees [choi2001multiscale, cheng2001multiscale] or neural networks [kim2009multiscale]. The main limitation of these approaches is that they rely on training data to describe the texture parameters of the images. Additionally, even though there exist works in the literature that attempt to consider coefficients at multiple scales for image segmentation, up to our knowledge, there is no method in the state of the art that exploits these scales to segment images at the pixel level based on a multifractal model.
@@
This work introduces a new multifractal image segmentation approach for the joint parameter and label estimation of heterogeneous images via a multiscale decomposition and a Bayesian framework. The proposed segmentation approach is fully unsupervised and is conducted in the wavelet leader domain, in which the scale-free textures are parameterized by the multifractality parameter. The contributions of this paper are the following:
@@
    First, a computationally efficient Fourier-based Whittle approximation for the multifractality parameter estimation on heterogeneous multifractal images is introduced. This contribution relies on the efficient model developed in [wendt2018multifractal] that is extended and adapted to irregular shapes for modeling the different piecewise textures of an image.
@@
    Third, a multiscale Potts Markov random field that considers the intra-scale and inter-scale dependencies in the wavelet leader domain is considered. This random field uses a multiscale graph structure that exploits spatial and scale image correlations. Moreover, the granularity coefficients of the proposed multiscale Potts-Markov random field are efficiently and jointly sampled with the labels.
@@
    Fourth, we conduct extensive numerical experiments and comparisons with state-of-the-art image segmentation methods for simulated and real-world images, showing the interest of the proposed framework.
@@ 2.1 Local Regularity and Multifractal Spectrum
MFA enables the study of local regularity of functions. To that end, it makes use of a quantity that measures the so-called Hölder exponent $h(\textbf{t})$ at a location t, and a global description of its spatial fluctuations, the multifractal spectrum $\mathcal{D}(h)$ [pesquet2002stochastic].
@@
Here, large (small) values of $h(\textbf{t}_{0})$ correspond to locations where $\mathbf{F}$ is locally smooth (irregular), respectively [pustelnik2016combining].
@@
The estimation of $\mathcal{D}(h)$ is the main target of MFA. However, it is unfeasible using its formal definition (1). Indeed, the estimation relies on a multiscale decomposition and the so-called multifractal formalism (for more details, see [wendt2007bootstrap, wendt2018multifractal, jaffard2004wavelet]). The current state-of-the-art multifractal formalism is based on the wavelet leaders of the two-dimensional (2D) discrete wavelet transform (DWT) coefficients of the given image and is recalled next.
@@ 2.2 Multifractal Formalism based on Wavelet Leaders
Wavelet Leaders.    The 2D orthonormal DWT of $\textbf{F}\in\mathbb{R}^{N\times N}$ is obtained by considering four 2D filters $\phi^{(m)}(\textbf{n})$, m=0,,3m=0,...,3, defined as the tensor products of a low-pass filter $\phi(n)$ (scaling function) and a high-pass filter $\psi(n)$ (wavelet function characterized by $N_{\psi}>0$ vanishing moments). Precisely, the four 2D filters are written as [combrexelle2015bayesian, wendt2018multifractal]:
@@
where $\psi^{(0)}(\textbf{n})$ is the 2D low-pass filter and $\{\psi^{(m)}(\textbf{n})\}_{m=1}^{3}$ are the 2D high-pass filters. Let $D^{(m)}_{\textbf{F}}(j,\mathbf{n})$ denotes the 2D DWT of F, with $D^{(0)}_{\textbf{F}}(0,\mathbf{n})\triangleq\textbf{F}$. The DWT coefficients $\big\{D_{\textbf{F}}^{(m)}(j,\mathbf{n})\big\}_{m=1}^{3}$, and the approximation coefficients $D_{\textbf{F}}^{(0)}(j,\mathbf{n})$, for $j\geq 1$, are obtained by iteratively convolving $\psi^{(m)}$ with the coefficients $D_{\textbf{F}}^{(0)}(j-1,\cdot)$, followed by decimation [wendt2018multifractal]. For MFA purposes, the 2D DWT coefficients are normalized as $d_{\textbf{F}}^{(m)}(j,\mathbf{n})\triangleq 2^{-j}D_{\textbf{F}}^{(m)}(j,\mathbf{n})$, for $m=1,2,3$.
The wavelet leaders are defined in [jaffard2006wavelet] as the largest DWT coefficient magnitudes within the spatial neighborhood $3\lambda_{j,\textbf{n}}$ over all finer scales:
[…]
where $\lambda_{j,\textbf{n}}$ is the dyadic cube of side length $2^{j}$ at position $2^{j}\textbf{n}$, i.e., $\lambda_{j,\textbf{n}}=\{[2^{j}n_{1},2^{j}(n_{1}+1)),[2^{j}n_{2},2^{j}(n_{2}+1))\}$ [lerner2019intuitive, Jaffard2015], and $3\lambda_{j,\textbf{n}}=\bigcup_{i_{1},i_{2}\in\{-1,0,1\}}\lambda_{j,(n_{1}+i_{1},n_{2}+i_{2})}$ (see [jaffard2006wavelet] for more details about wavelet leaders and related bounds).
It has been shown that the coefficients $c_{p}$ are tied to the wavelet leaders $L_{j,\mathbf{n}}$ via the key relation [combrexelle2015bayesian]
@@
for $c_{2}<0$, see [wendt2018multifractal, combrexelle2015bayesian] for further details. These so-called log-cumulants provide a relevant summary of $\mathcal{D}(h)$, e.g., $c_{1}$ is the mode of $\mathcal{D}(h)$ that can be interpreted as the average smoothness of the image $\mathbf{F}$ and is related to the Hurst parameter [mandelbrot1968fractional], $c_{2}$ is related to the width of $\mathcal{D}(h)$ and captures the degree of fluctuations of the local regularity of $\mathbf{F}$ and is the most important feature of multifractal functions, termed multifractality parameter. Figure 1 presents a schematic illustration of the theoretical multifractal spectrum for two homogeneous 2D multifractal random walks, shown for $c_{1}=0.5,c_{2}=-0.02$ (blue curve) and $c_{1}=0.98,c_{2}=-0.08$ (black curve) [wendt2007bootstrap, wendt2007multifractality, Wendt20091100] (note that these are the multifractal spectra of the composed image of Fig. 4-a-top).
Figure 1: Multifractal spectrum $\mathcal{D}(h)$ schematic representation from two 2D multifractal random walk images, where $c_{1}$ is the mode and $c_{2}$, the width, is the multifractality parameter; $(c_{1},c_{2})$ capture most of the local regularity information of the image.
@@
[…]
@@ 2.3 Estimation of $c_{2}$ using a Fourier-based Bayesian model
Let ${\ell}_{j,\textbf{n}}=\ln(L_{{j,\textbf{n}}})$ denote the centered logarithms of the wavelet leader coefficients, referred to as log-leaders, at scale $j$ and at position n. The vector of log-leaders, $\boldsymbol{\ell}_{{j}}\in\mathbb{R}^{N_{j}^{2}}$, is obtained by vectorization using the lexicographic order, where $N_{j}=N/2^{j}$. Note that $j\in\llbracket j_{1},j_{2}\rrbracket$, $1\leq j_{1}\leq...\leq j_{2}\leq J$, in which $j=1$ and $j=J$ index the finest and the coarsest scales, respectively, where $\llbracket a,b\rrbracket$ denotes the set of integers ranging from $a$ to $b$. Further, let $\boldsymbol{\ell}\triangleq[\boldsymbol{\ell}_{j_{1}}^{T},...,\boldsymbol{\ell}_{j_{2}}^{T}]^{T}$ be the vector obtained by concatenation of the log-leaders at scales $j_{1},...,j_{2}$. According to [wendt2018multifractal, combrexelle2015bayesian], the distribution of the log-leaders $\boldsymbol{\ell}_{j}$ can be well-approximated by a multivariate Gaussian distribution with density given by:
$p(\boldsymbol{\ell}_{j}|\theta_{1},\theta_{2})\propto{{\big|\,\,\mathbf{\Sigma}_{j}(\boldsymbol{\theta})\big|^{-1/2}}}{\exp\Big(-\frac{1}{2}\boldsymbol{\ell}_{j}^{T}\mathbf{\Sigma}_{j}(\boldsymbol{\theta},\left\lVert\Delta\textbf{n}\right\rVert)^{-1}\boldsymbol{\ell}_{j}\Big)}.$ (6)
Assuming independence between scales, one obtains:
@@
where $|\textbf{A}|$ denotes the determinant of a matrix A, $\boldsymbol{\theta}\triangleq[{\theta}_{1},{\theta}_{2}]^{T}$, $c_{2}={\theta}_{1}$, ${\theta}_{2}$ is related to the model adjustment constant $c_{2}^{0}$, and $\mathbf{\Sigma}_{j}(\boldsymbol{\theta},\left\lVert\Delta\textbf{n}\right\rVert)\approx\textrm{Cov}(\boldsymbol{\ell}_{j,\mathbf{n}},\boldsymbol{\ell}_{j,\mathbf{n}+\Delta\mathbf{n}})$ is a covariance model [combrexelle2015bayesian], defined by the radial symmetric function
@@
where $r\triangleq\left\lVert\Delta\textbf{n}\right\rVert$, $g_{1}$ and $g_{2}$ are functions defined as $g_{1}(j,r)=\max\{0,-\ln\big((r+1)/(r_{j}+1)\big)\}$ and $g_{2}(j,r)=\max\{0,-\ln(r+1)/\ln 4\}$ with $r_{j}=\lfloor N_{j}/4\rfloor$ [leon2022bayesian].
The likelihood (7) is difficult to be used in a numerical algorithm but can be efficiently approximated in the spectral domain by formulating an equivalent Fourier-domain data augmented model introduced in [wendt2018multifractal]. This statistical model is based on a Whittle approximation and data augmentation allowing the suitable design of conjugate priors for the parameters ${\theta}_{1},{\theta}_{2}$ [wendt2018multifractal, combrexelle2015bayesian]. More precisely, (7) is approximated as
@@
$p\big(\textbf{x},\boldsymbol{\mu}|{\theta}_{1},{\theta}_{2}\big)\propto{\theta}_{1}^{-S}\exp\bigg(-\frac{1}{{\theta}_{1}}(\textbf{x}-\boldsymbol{\mu})^{H}G^{-1}_{1,s}(\textbf{x}-\boldsymbol{\mu})\bigg){\theta}_{2}^{-S}\exp\bigg(-\frac{1}{{\theta}_{2}}\boldsymbol{\mu}^{H}G^{-1}_{2,s}\boldsymbol{\mu}\bigg),$ (10)
[…]
@@
where $s\in\big\{(j,\textbf{m}):j=\llbracket j_{1},j_{2}\rrbracket,\textbf{m}\in\Omega^{\dagger}_{j}\big\}$, $\Omega^{\dagger}_{j}$ indexes low frequencies (see [wendt2018multifractal, leon2022bayesian] for further details), and $\boldsymbol{\omega}_{s}=2\pi\mathbf{m}/\sqrt{N_{j}}$ is the given Fourier frequency, with $(j,\textbf{m})\in s$. Moreover, the matrices
${G}_{i,s}\triangleq\sum_{\textbf{n}\in\llbracket-{N_{j}},{N_{j}}\rrbracket^{2}}g_{i}(j,\left\lVert\textbf{n}\right\rVert)\exp{(-i\,\textbf{n}^{T}\boldsymbol{\omega}_{\mathbf{s}})},$ (12)
for $i=1,2$ are real and strictly positive matrices defined on $s$. The corresponding spectral covariance of the asymptotic approximation (10) is expressed as
@@
which leads to standard conditional distributions and fast estimation when used in a Bayesian model [wendt2018multifractal, leon2022bayesian].